OpenStudy (anonymous):

Find all real values of x such that f(x) = 0. f(x) = x^2 − 8x + 15 show steps please

OpenStudy (anonymous):

f(x)=0 f(x)=x^2-8x+15 0=x^2-8x+15 0=(x-5)(x-3) x-5=0 x-3=0 x=5 x=3